A random sample of 50 students is chosen from a large population whose diastolic blood pressures has a standard deviation of 5mm Hg. If the 50 students gave a mean pressure of 80 mm Hg, compute the 88.12% confidence interval of the mean of the diastolic pressures of all students.


1
Expert's answer
2021-05-18T06:56:22-0400

The following information is provided from 88.12% confidence interval for population mean "\mu" .

Sample mean "\overline{X}=80"

Sample standard deviation "\:\:\left(s\right)=5\:"

Sample size "\:\left(n\right)=50"


The critical value for "\u03b1=0.119" and "df=n\u22121=49" degrees of freedom is "\:\:\:t_c=Z_{1-\frac{\alpha }{2};n-1}=1.588". The corresponding confidence interval is computed as shown below:

"CI=\left(\overline{X}-t_c\times \frac{s}{\sqrt{n}},\:\overline{X}+t_c\times \:\frac{s}{\sqrt{n}}\:\right)"

"CI\:\:=\left(80-1.588\times \:\:\frac{5}{\sqrt{50}},\:80+1.588\times \:\:\:\frac{5}{\sqrt{50}}\:\right)"

"CI=(78.877, 81.123)"


Therefore, based on the data provided, the 88.12% confidence interval for the population mean is 78.877<μ<81.123, which indicates that we are 88.12% confident that the true population mean μ is contained by the interval (78.877,81.123).